OpenStudy (anonymous):

Find the common difference of the arithmetic sequence. 9, 13, 17, 21, . .

OpenStudy (anonymous):

I think the answer is 4 because the series is 9,13,17,21.... 9+4=13 13+4 = 17 17 + 4 = 21 And also 13-9 = 4 17 - 13 = 4 21-17 = 4 so the answer is 4

OpenStudy (anonymous):

So, we have: \[ 13-9=17-13=21-17=\;? \]What would that be?